# Mutatis Mutandis: Revisiting the Comparator in Discrimination Testing

### v2

@@ 5.2. Discrimination Discovery
Discrimination discovery tools (Kamiran et al., 2013; Ruggieri et al., 2010; Pedreschi et al., 2008) represent the first attempt at what we now call algorithmic fairness. These works took the standard methods and combined them with data mining methods, shifting the focus from generating or finding the experimental data to exploring a data produced by a decision process. In part, this shift in focus was field-driven: within data mining and overall knowledge discovery, the focus of the problem formulation is on developing new and efficient algorithms for extracting information from large datasets. This is a different focus from the prevalent inferential statics approach in which the creation of the data is as important as using the most efficient model. It came with a change in how we test for discrimination, especially with the adoption of modern decision flows, as the attention was now on a consistent, algorithmic decision-maker. Hence, to test whether the decision-maker discriminated, we could now look at the data it produced.
Under these tools, the setup in Section 2.1 started to become standard. It was a natural transition. Both the decision-maker in question and the tools for evaluating it were scalable. These works borrowed considerably both from non-discrimination law and from the standard methods. Works by Pedreschi et al. (2008) and Ruggieri et al. (2010) provided the first formalization to equal opportunity (and, implicitly, to equalized odds) considerably before Hardt et al. (2016). Kamiran and Calders (2009) did the same for demographic parity. Thanh et al. (2011) is an illustrative example of the discrimination discovery wave, proposing a $k$-NN implementation of situation testing.
These tools focus on defining the best distance to reach the comparator for a given complainant. Some of them are closer to the MM comparator than to the CP comparator, though. Thanh et al. (2011) clearly use the CP comparator as they minimize the distance between the complainant and potential comparators based only on the set of non-protected attributes. Zhang et al. (2016), instead, offer an extension to (Thanh et al., 2011) via causal weights, focusing on defining a distance function that prioritizes the important non-protected attributes when finding the comparator for a complainant. Similarly, Qureshi et al. (2020) use a propensity score weight for the distance function with the protected attribute as the treatment. None of these tools, however, explicitly address the problem with using the CP comparator or the need to derive the MM comparator. There is no discussion on how the protected attribute might influence all other attributes.
@@ 5.3. Algorithmic Fairness
In the past decade, discrimination discovery took a broader, more multidisciplinary focus and became algorithmic fairness. The objectives are the same, but algorithmic fairness is broader in scope, usually aiming to reformulate standard ML problems under a fairness objective (Barocas et al., 2019). In turn, many of these tools, even if they are not intended to prove discrimination, are useful for tracking potentially discriminatory decision-making. Given the counterfactual model of discrimination and our focus on the comparator, the main fairness works useful for discrimination are counterfactual fairness (Kusner et al., 2017) and individual fairness (Dwork et al., 2012). The former establishes the same treatment of the factual and its counterfactual instance under the same decision; the latter establishes the same treatment of similar instances under the same decision. Under the counterfactual model of discrimination, these two concepts intersect as we use the similarity between individuals to approximate the comparison between an individual and its counterfactual-self. Works like the FlipTest (Black et al., 2020) or counterfactual situation testing (Álvarez and Ruggieri, 2023) build upon these two definitions. Overall, this holds for other similar works including (Dutta et al., 2021; Anthis and Veitch, 2023; Salimi et al., 2019).
@@
We view algorithmic fairness tools based on generative models as central to implementing the MM comparator. It includes that are not necessarily concerned with discrimination testing but on, e.g., learning fair representations (Zemel et al., 2013). Tools that are more focused on proving discrimination, meaning along the same lines as discrimination discovery, still face the key challenge of linking unfairness to discrimination (Weerts et al., 2023). These tools are useful, but it remains to be seen how they fit into the pipeline for establishing discrimination. This is not a technical challenge in-itself, but one on whether the technology is embraced or not by the relevant stakeholders. Another key challenge, which links directly to the MM comparator, is how these tools address indirect discrimination as most algorithmic decision-makers are not allowed to use the protected attribute as input. As illustrated in Section 4, the choice of comparator affects the number of discrimination cases identified. These tools, along with the discrimination discovery ones, are capable of generating more complex comparators than the standard tools. The question then is whether they should aim or not at implementing the MM comparator over the CP comparator. Our view is, at a minimum, they should allow for the possibility of choosing between the two.
@@ 6. Conclusion
We discussed extensively the problem of discrimination testing by focusing on the notion of comparator. We presented the derivation of the comparator as a causal modeling problem, emphasizing the role ML methods like generative models can play for better discrimination testing methods. We introduced the ceteris paribus (CP) comparator and mutatis mutandis (MM) comparator distinction; the former refers to the standard comparator and the latter represents a new kind of comparator that departs from idealized comparisons. Experiments showcase the impact this distinction has on discrimination testing and motivates the need to consider, at least, the option of being able to choose the kind of comparator when implementing a tool for discrimination testing.
@@
This work comes from a ML perspective on discrimination testing. We have done our best in incorporating other fields and their views on the need for the MM comparator. Future work should revisit the problem of discrimination testing by considering the full legal procedure for establishing discrimination. Future work should consider other generative models for implementing the MM comparator beyond SCM. We are certain that the MM comparator and its departure from the CP comparator, both as a conceptual and modeling problem, will motivate further fair ML work on discrimination testing.
@@ Acknowledgements.
This work has received funding from the European Union’s Horizon 2020 research and innovation program for the MSCA project NoBIAS (g.a. No. 860630) and for the project FINDHR (g.a. No. 101070212). This work reflects only the authors’ views and the European Research Executive Agency is not responsible for any use that may be made of the information it contains.

### v3

@@ 5.2. Discrimination Discovery
The discrimination discovery tools (Kamiran et al., 2013; Ruggieri et al., 2010; Pedreschi et al., 2008) represent the first attempt at what we now call algorithmic fairness. These works took the standard methods of the previous section and combined them with data mining methods, shifting the focus from generating or finding the experimental data to exploring data already produced by a decision process. This shift in focus was in part field-driven: within data mining and overall knowledge discovery, the focus of the problem formulation is on developing new and efficient algorithms for extracting information from large datasets. It is a different focus from the prevalent inferential statistics approach in which the creation of the data is as important as using the most efficient model. But this shift also came with a change in how we test for discrimination, especially with the adoption of modern decision flows, as the attention is now on a consistent, algorithmic decision-maker. Hence, to test whether the decision-maker discriminated, we could now look at the data it produced.
The setup in Section 2.1 started to become standard under these tools. It was a natural transition from the standard tools as these works borrowed considerably both from non-discrimination law and from the standard methods. For example, works by Pedreschi et al. (2008) and Ruggieri et al. (2010) provided the first formalization to equal opportunity (and, implicitly, to equalized odds) considerably before Hardt et al. (2016). Kamiran and Calders (2009) did the same for demographic parity. Thanh et al. (2011) is an illustrative example of the discrimination discovery wave, proposing a $k$-NN implementation of situation testing (Rorive, 2009).
These tools largely focus on defining the best distance to reach the comparator for a given complainant.
Some of the tools are closer to the MM comparator, but most of them implement the CP comparator. Thanh et al. (2011) clearly use the CP comparator as they minimize the distance between the complainant and potential comparators based only on the set of non-protected attributes. Zhang et al. (2016), instead, offer an extension to (Thanh et al., 2011) via causal weights, focusing on defining a distance function that prioritizes the important non-protected attributes when finding the comparator for a complainant. Similarly, Qureshi et al. (2020) use a propensity score weight for the distance function with the protected attribute as the treatment. None of these tools, however, explicitly address the problem with using the CP comparator or the need to derive the MM comparator. For example, there is no discussion on how the protected attribute might influence all other attributes.
@@ 5.3. Algorithmic Fairness
Discrimination discovery took a broader, more multidisciplinary focus over the past decade and became algorithmic fairness. The objectives are the same, but algorithmic fairness is broader in scope, usually aiming to reformulate standard ML problems under a fairness objective (Barocas et al., 2019). In turn, many of these tools, even if they are not intended to prove discrimination, are useful for tracking potentially discriminatory decision-making. Unsurprisingly, the terms unfair and discriminatory are used interchangeably with algorithmic fairness, sometimes incorrectly (Weerts et al., 2023). As discussed in Section 2.2, discrimination claims must align with non-discrimination law in terms of the protected attribute and testing tools represent one aspect of establishing discrimination.
Given the counterfactual model of discrimination and our focus on the comparator, the main algorithmic fairness works that are central to discrimination testing are counterfactual fairness (Kusner et al., 2017) and individual fairness (Dwork et al., 2012). The former establishes the same treatment of the factual and its counterfactual instance under the same decision; the latter establishes the same treatment of similar instances under the same decision. Both fairness definitions allude explicitly to the complainant-comparator pair. Under the counterfactual model of discrimination, these two concepts intersect as we use the similarity between individuals to approximate the comparison between an individual and its counterfactual-self. Works like the FlipTest (Black et al., 2020) or CST (Álvarez and Ruggieri, 2023) build upon these two definitions. Overall, this holds for other similar works including (Dutta et al., 2021; Anthis and Veitch, 2023; Salimi et al., 2019). Now algorithmic fairness is a vast and ongoing field, with research on fairness definitions and their links to non-discrimination law being one active branch (Wachter et al., 2020, 2021). Moving forward, it is important that the CP versus MM comparator discussion is present when looking at (algorithmic) discrimination. The choice of method or definition matters beyond providing a complete view when running experiments. Such a choice, as discussed, has wider consequences on what we consider discriminatory that are normative concerns rather than modeling concerns. Each method, as also argued by Wachter et al. (2020), operationalizes a specific ML “intervention” to the systematic unfairness behind the discrimination claim. Of this modeling choice we must be aware.
Moving forward, especially given our formulation of $\tilde{\mathbf{X}}$ in Section 3.3, we view algorithmic fairness tools based on generative models as central to the MM comparator. In particular, methods concerned with learning fair representations. This research direction started with Zemel et al. (2013) and has led to multiple proposed methods over the past decade (e.g., Cerrato et al. (2023, 2022); Ganin et al. (2016); Louizos et al. (2016); Moyer et al. (2018); Xie et al. (2017); Madras et al. (2018)). The goal with these methods is to learn a fair representation of the individual, meaning learning a representation of seemingly neutral attributes that removes information from the protected attribute while keeping valuable information of the attributes in question useful for a downstream task. We recommend Cerrato et al. (2024) for a survey on fair representations. Since those complex models tackle the problem (5), they could be, in principle, repurposed for discrimination testing with the goal of deriving an MM comparator. Future research is open along this interesting line. Our approach relies on SCM, for which such a derivation boils down to counterfactual generation (cfr. Section 1.1).
@@
Algorithmic fairness tools that are more focused on proving discrimination, meaning along the same lines as discrimination discovery, still face the key challenge of linking unfairness to discrimination (Weerts et al., 2023). These tools are useful, but it remains to be seen how they fit into the pipeline for establishing discrimination. This is not a technical challenge in-itself, but one on whether the technology is embraced or not by the relevant stakeholders. Another key challenge, which links directly to the MM comparator, is how these tools address indirect discrimination as most algorithmic decision-makers are not allowed to use the protected attribute as input. As illustrated in Section 4, the choice of comparator affects the number of discrimination cases identified. These tools, along with the discrimination discovery ones, are capable of generating more complex comparators than the standard tools. The question then is whether they should aim or not at implementing the MM comparator over the CP comparator. Our view is, at a minimum, they should allow for the possibility of choosing between the two.
@@ 6. Conclusion
This work discussed the problem of discrimination testing by focusing on the comparator. First, we presented the derivation of the comparator as a causal modeling problem. It is not the most popular view on discrimination testing (Kohler-Hausmann, 2019), but an important one as it opens impactful opportunities for ML models. Second, we introduced the ceteris paribus (CP) comparator and mutatis mutandis (MM) comparator distinction. The former refers to the standard comparator and the latter represents a new kind of comparator that departs from idealized comparisons. This distinction is novel and will help not only in classifying existing methods but also to better inform new ones and how their choice of similarity between individuals unavoidably impacts normative concerns related to discrimination. Additionally, the illustrative experiment based on situation testing (Thanh et al., 2011) and counterfactual situation testing (Álvarez and Ruggieri, 2023, 2025) showcased the impact this distinction has on discrimination testing, motivating the need to consider, at least, the option to choose the kind of comparator when implementing a tool for discrimination testing.
We focused on causal modeling, particularly, structural causal models (SCM) for generating the MM comparator. We stress that said comparator is not tied to this modeling choice, meaning other ML approaches are possible and encouraged. Fair representation learning (Cerrato et al., 2024) is the most appealing direction in terms of generative ML for the MM comparator. That said, given our causal modeling choice for generating the MM comparator, two limitations are clear and worth discussing. First, we use the same procedure as in Kusner et al. (2017) to generate the suitable imaginary comparators, meaning using Pearl et al. (2016)’s abduction, action, and prediction steps. It remains a common procedure that can rely on either a frequentist or Bayesian approach. However, it is also a procedure that requires sufficient knowledge about the SCM generating the data, which is complex and potentially non-scalable in practice. Other causal ML approaches, such as causal normalizing flows (Javaloy et al., 2023), offer an alternative to generating the counterfactual distribution without having to be explicit about model specifications in the SCM. We do not explore these methods as our focus is on the CP versus MM comparison. We leave this as future work. Second, our extension of the standard discrimination testing setup (Section 3.3) centers on generating an adjusted representation of the non-protected attributes. In that sense, it aligns well with the problem of fair representation learning. However, an alternative approach could be to define an adjusted similarity function with the same objective. We do not consider this approach as, given our use of SCM and overall emphasis on generative ML, the generation of $\tilde{\mathbf{X}}$ fits better. We too leave this as future work. Exploring different kinds of data-driven distance functions (Blanco-Mallo et al., 2023) as well as searching for similar individuals based on their latent space representation are potential directions to explore.
Overall, defining a distance function can be as challenging as agreeing on a SCM for generating the MM comparator in the form of $\tilde{\mathbf{X}}$. Consider, for instance, the discussions on implementing individual fairness by Dwork et al. (2012), which always comes back to the circularity argument in Section 2 inspired by Westen (1982). Defining a similarity measure between individuals is inherently contentious. At least with the SCM our answer to the what would have been question is answered by the counterfactual distribution, in a way settling the “choice of similarity metric” debate as a generative question. Redirecting the focus on a distance function, we believe, would still require to define an imagined objective to learn the distance. In fact, said distance could be based on evaluating the factual and counterfactual distributions. This is a promising line of research.
Further, given now the proposed CP and MM comparator distinction, a clear direction for future work is to move beyond an illustrative experiment and survey suitable methods for the MM comparator in a robust experimental setting. The goal here was to motivate and formalize such distinction, which is by all means not trivial and often not shared among all of those in the discrimination testing field. Future work should move beyond SCM and consider the methods already mentioned as well as other fair ML benchmark datasets to explore the potential limitations of the MM comparator as a modeling problem.
@@
Finally, this work comes from a ML perspective on discrimination testing. We have done our best in incorporating other fields and their views on the need for the MM comparator. In doing so, we focused on one aspect—that of providing prima facie evidence—of the full pipeline for establishing discrimination. We hope future legal work complements what we have presented here as already the choice of a MM over a CP comparator would expand the current pipeline, requiring additional legal commentary. We are certain that the MM comparator and its departure from the CP comparator, both as a conceptual and modeling problem, will motivate further fair ML work on discrimination testing. As shown in Example 1.1, it is not straightforward to determine who Clara should be compared to when evaluating her discrimination claim. The introduction of the MM comparator helps facilitate this complex and important discussion.
@@ Acknowledgements.
We would like to thank Vijay Keswani, Otto Sahlgren, Mattia Cerrato, Alesia Vallenas Coronel, Marcello Di Bello, Nicolò Cangiotti, Francesco Nappo, and Michele Loi for their feedback at different stages of this work. This work has received funding from the European Union’s Horizon 2020 research and innovation program for the MSCA project NoBIAS (g.a. No. 860630) and for the project FINDHR (g.a. No. 101070212). It reflects only the authors’ views and the European Research Executive Agency is not responsible for any use that may be made of the information it contains.
Appendix A On The Trouble with Disparity
Appendix B Supplementary Material for Counterfactual Situation Testing
In this section, we present additional technical material for CST (Álvarez and Ruggieri, 2023, 2025). CST relies on the k-NN algorithm, building the neighborhoods of the control and test groups using a distance function $d(\cdot,\cdot)$ between two individual profiles or tuples. Such function only looks at non-protected attributes. Let us define the between tuple distance $d(x_{1},x_{2})$ as:
(8) $d(x_{1},x_{2})=\frac{\sum_{i=1}^{|X|}d_{i}(x_{1,i},x_{2,i})}{|X|}$
such that $d(x_{1},x_{2})$ averages the sum of the per-attribute distances $d_{i}(x_{1,i},x_{2,i})$ across all attributes in $X$. CST handles non-normalized attributes but, as default, we normalize them to insure comparable per-attribute distances.
The $d_{i}$ used depends on the type of the i-th attribute. It equals the overlap measurement ($ol$) if the attribute $X_{i}$ is categorical; otherwise, it equals the normalized Manhattan distance ($md$) if the attribute $X_{i}$ is continuous, ordinal, or interval. Hence, $d(\cdot,\cdot)$ amounts to Gower’s distance (Gower, 1971). We recall $md$ and $ol$ distances below:
(9) $md(x_{1,i},x_{2,i})=\frac{|x_{1,i}-x_{2,i}|}{(\max(X_{i})-\min(X_{i}))}$
(10) $ol(x_{1,i},x_{2,i})=\begin{cases}1&\text{if }x_{1,i}\neq x_{2,i}\\ 0&\text{otherwise}.\end{cases}$
The main objective of CST is to estimate the difference in proportion of negative decision outcomes for each complainant’s control and test groups:
(11) $\Delta p=p_{c}-p_{t}$
with $p_{c}$ and $p_{t}$ represents the count of individuals with a negative decision outcome, respectively, in the control group and test group. Recall, that the control group is centered on the complainant $i$ and the test group on its (generated counterfactual) comparator $i^{\prime}$.
For our purposes, there is potential individual discrimination toward a complainant if $\Delta p>\tau$, where $\tau\in[-1,1]$ represents some accepted deviation. Often, $\tau=0$. This would amount to a general prima facie discrimination claim. CST can also equip such claim with confidence intervals, addressing concerns on whether it is statistically significant or not. For more details, we invite the reader to see Álvarez and Ruggieri (2025).